# Multimodal Industrial Anomaly Detection by Crossmodal Reverse Distillation

### v2

@@ 4.2 Crossmodal Assistants
Method/Category Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Avg.
DepthGAN [7] 53.8/42.1 37.2/42.2 58.0/77.8 60.3/69.6 43.0/49.4 53.4/25.2 64.2/28.5 60.1/36.2 44.3/40.2 57.7/63.1 53.2/47.4
DepthAE [7] 64.8/43.2 50.2/15.8 65.0/80.8 48.8/49.1 80.5/84.1 52.2/40.6 71.2/26.2 52.9/21.6 54.0/71.6 55.2/47.8 59.5/48.1
DepthVM [7] 51.3/38.8 55.1/32.1 47.7/19.4 58.1/57.0 61.7/40.8 71.6/28.2 45.0/24.4 42.1/34.9 59.8/26.8 62.3/33.1 55.5/33.5
BTF [20] 91.8/97.6 74.8/96.9 96.7/97.9 88.3/97.3 93.2/93.3 58.2/88.8 89.6/97.5 91.2/98.1 92.1/95.0 88.6/97.1 86.5/95.9
M3DM [31] 99.4/97.0 90.9/97.1 97.2/97.9 97.6/95.0 96.0/94.1 94.2/93.2 97.3/97.7 89.9/97.1 97.2/97.1 85.0/97.5 94.5/96.4
Shape-guided [11] 98.6/98.1 89.4/97.3 98.3/98.2 99.1/97.1 97.6/96.2 85.7/97.8 99.0/98.1 96.5/98.3 96.0/97.4 86.9/97.5 94.7/97.6
AST [26] 98.3/97.0 87.3/94.7 97.6/98.1 97.1/93.9 93.2/91.3 88.5/90.6 97.4/97.9 98.1/98.2 100/88.9 79.7/94.0 93.7/94.4
MMRD [15] 99.9/98.6 94.3/99.0 96.4/99.1 94.3/95.1 99.2/99.0 91.2/90.1 94.9/99.0 90.1/99.0 99.4/98.7 90.1/98.2 95.0/97.6
CFM [12] 98.8/98.0 87.5/96.6 98.4/98.2 99.2/94.7 99.7/95.9 92.4/96.7 96.4/98.2 94.9/98.3 97.9/97.6 95.0/98.2 96.0/97.2
CRD (Ours) 99.6/97.7 96.1/98.1 98.2/99.1 98.5/98.2 99.7/98.7 90.1/91.0 99.2/99.4 97.6/99.5 100/98.2 83.3/97.4 96.2/97.7
Table 1: Anomaly detection and localization results I-AUROC/PRO (%) on MVTec 3D-AD with the best in bold and the second-best underlined.
4.3 Training Objectives
@@
Secondly, to ensure the effectiveness of Crossmodal Filter, it is crucial to align the mapping features $F_{BP}^{3D/2D}$ within these filters to the encoder features. The training losses of 2D and 3D branches are calculated as:
@@
Finally, Crossmodal Amplifier needs to be optimized,, which includes the optimization of feature mapping and feature fusion parameters $\{w_{1},w_{2}\}$. The total loss function is
@@
The overall loss $\mathcal{L}_{CRD}$ is expressed as
$\mathcal{L}_{CRD}=(\mathcal{L}^{2D}_{D}+\mathcal{L}^{2D}_{CF}+\mathcal{L}^{2D}_{CA})+(\mathcal{L}^{3D}_{D}+\mathcal{L}^{3D}_{CF}+\mathcal{L}^{3D}_{CA})$ (9)
Since each part has independent optimizers and inputs, the hyper-parameters are set to 1.
4.4 Anomaly Scoring
@@
Methods Candy Cane Chocolate Cookie Chocolate Praline Confetto Gummy Bear Hazelnut Truffle Licorice Sandwich Lollipop Marsh- mallow Peppermint Candy Avg.
M3DM [31] 62.4/90.6 95.8/92.3 95.8/80.3 100/98.3 88.6/85.5 75.8/68.8 94.9/88.0 83.6/90.6 100/96.6 100/95.5 89.7/88.2
AST [26] 57.4/51.4 74.7/83.5 74.7/71.4 88.9/90.5 59.6/58.7 61.7/59.0 81.6/73.6 84.1/76.9 98.7/91.8 98.7/87.8 78.0/74.4
MMRD [15] 85.4/97.5 100/97.0 94.6/94.2 99.8/98.5 90.8/91.7 74.7/68.0 96.6/97.0 98.4/94.1 100/99.0 100/99.2 94.0/93.6
CFM [12] 68.0/94.2 93.1/90.2 95.2/83.1 88.0/96.5 86.5/87.5 78.2/76.2 91.7/79.1 84.0/91.3 99.8/93.9 96.2/94.9 88.1/88.7
CRD (Ours) 86.7/98.0 100/94.1 98.1/96.8 98.1/98.1 87.2/93.6 72.8/77.3 97.1/98.1 93.0/94.4 100/98.9 99.5/98.1 93.3/94.7
Table 2: Anomaly detection and localization results I-AUC/PRO (%) on Eyecandies with the best in bold and the second-best underlined.
5 Experiments
5.1 Experimental Settings
@@
Dataset. We conduct experiments MVTec 3D-AD [7] and Eyecandies [9]. MVTec 3D-AD is a 3D industrial anomaly detection dataset featuring 10 categories of industrial objects. The training set consists of 2,656 normal samples, each including an RGB image and corresponding 3D scan information, while the validation set contains 294 normal samples. The test set comprises 1,197 samples, including 948 anomalous samples and 249 normal samples, with pixel-level annotations provided for evaluation. For our experiments, we use the processed depth maps derived from the 3D scans as the 3D modality. Eyecandies is a synthetic dataset of 10 candy categories in an industrial conveyor scenario, providing RGB images, depth maps, and surface normal maps. It contains 10k training samples, 1k validation samples, and 4k test samples. In our experiments, we select normal maps as the 3D modality and chose images under consistent lighting conditions as the 2D modality.
@@
Evaluation Metrics. The evaluation metrics for anomaly detection include image-level area under the receiver operating characteristic curve (I-AUC) and average precision (I-AP). For anomaly localization, pixel-level AUC (P-AUC) and AP (P-AP) are used. We also employ the per-region overlap (PRO) metric calculated according to the official code [7] for anomaly localization, which treats anomaly regions of different size equally.
@@
Implementation Details. All RGB images, depth maps and normal maps are scaled to $256\times 256$. The training is performed on a single Nvidia GTX 3090 GPU, and a separate detection model is trained for each category. The network backbone is WideResNet-50 [33]. The batch size is 16, and the training lasts for 200 epochs. Each branch is optimized using a separate Adam optimizer with a learning rate of 0.005. In inference, the anomaly maps of both branches are smoothed using a Gaussian filter with $\sigma=4$.
Figure 5: Qualitative ablation results of crossmodal assistants on MVTec 3D-AD.
5.2 Main Results
@@
Results on MVTec 3D-AD. Tab. 1 shows the quantitative image-level anomaly detection results and pixel-level anomaly localization results on the MVTec 3D-AD dataset. We compare our method CRD with several state-of-the-art 3D anomaly detection methods. CRD consistently outperforms all other methods in both I-AUC and PRO metrics, achieving an average I-AUC of 96.2% and an average PRO of 97.7%. Notably, CRD demonstrates superior performance in detecting and localizing anomalies across a wide range of categories, including some challenging categories such as Cookie and Peach. A subset of qualitative results on MVTec 3D-AD is provided in Fig. 1, showing that our proposed CRD accurately localizes the anomaly regions.
Results on Eyecandies. We evaluate our method on the Eyecandies dataset, as shown in Tab. 2. Our method CRD significantly outperforms all non-KD-based multimodal anomaly detection methods. While it slightly lags behind MMRD [15] which uses both depth maps and surface normal maps as 3D modalities in anomaly detection metrics I-AUC, it consistently achieves superior performance 94.7% in anomaly localization metrics PRO.
Method MVTec 3D-AD (%) Eyecandies (%) Avg. (%) Frame Rate (fps)
@@
CRD (Ours) 96.2 99.4 97.7 93.3 98.5 94.7 96.6 21.7
Table 3: Comparison of feature learning-based methods on MVTec 3D-AD and Eyecandies datasets under the same hardware setup. * indicates that MMRD has no reproducible code.
Comparison with Feature Learning-based Methods. We further compare CRD with feature learning-based methods including AST[26], MMRD[15], and CFM [12] on MVTec 3D-AD and Eyecandies. As in Table 3, CRD achieves the best average performance 96.6% and the highest frame rate 21.7 fps. Despite the multi-branch design, CRD remains lightweight due to its CNN-based architecture, outperforming other feature learning-based methods in efficiency.
5.3 Ablation Studies
@@
2D+3D+Crossmodal 96.2 99.4 97.7 98.8 45.5 87.5
Table 4: Ablation study results (%) of Multi-branch Distillation on MVTec 3D-AD.
@@
Table 5: Ablation study results (%) of Crossmodal Assistants on MVTec 3D-AD.
Study on Multi-branch Distillation. We conduct an ablation study in Tab. 4, which demonstrates that the multi-branch setup, combining both 2D and 3D modalities, significantly outperforms individual modality branches. The I-AUC metric of 93.9% for using both branches is higher than 86.8% for 2D branch only and 87.1% for 3D branch only, highlighting the complementary benefits of modality fusion.
Study on Crossmodal Assistants. We perform an ablation study to evaluate the impact of crossmodal assistants on anomaly detection performance, as shown in Tab. 5. The results reveal that both CF and CA contribute to improved performance, with the combination of both achieving the best results. Specifically, adding CF to the baseline improves the I-AUC from 93.9% to 94.2%, and incorporating CA also boosts the performance, achieving an I-AUC of 95.6%. The highest performance is achieved when both CF and CA are used together, resulting in an I-AUC of 96.2%.
We also provide visualizations of anomaly maps before and after incorporating crossmodal assistants, as in Fig. 5. Without crossmodal assistants, anomalies that are less prominent in a specific modality goes undetected in the corresponding branch. However, after introducing crossmodal assistants, this issue is significantly mitigated, resulting in a noticeable reduction in the anomaly miss rate across all modality branches.
@@
$16\times 16$ 93.6 99.3 97.4 97.9 45.2 86.7
$8\times 8$ 94.2 99.3 97.5 98.2 45.8 87.0
$4\times 4$ 93.8 99.3 97.4 98.0 45.3 86.8
Table 6: Ablation study results (%) of different downsampling sizes in Bottlneck Projection on MVTec 3D-AD.
Study on Bottlneck Projection in Crossmodal Filter. We investigate the effect of different downsampling sizes on the performance of the Crossmodal Filter, as shown in Tab. 6. When Bottleneck Projection reduces the feature size to $8\times 8$, the model achieves the best performance across multiple metrics. It is observed that excessively large downsampling size (e.g., $4\times 4$) leads to information loss, while small downsampling size (e.g., $16\times 16$) makes it difficult to filter out abnormal interference effectively. Therefore, we conclude that selecting an appropriate downsampling size is crucial for maintaining a balance between information preservation and interference removal.

### v3

@@ IV-B2 Crossmodal Amplifier
IV-C Training Objectives
@@
Secondly, to ensure the effectiveness of Crossmodal Filter, it is crucial to align the mapping features $F_{BP}^{3D/2D}$ output by Crossmodal Filters to the encoder features, as shown in Fig. 4. The training losses of 2D and 3D branches are calculated as
@@
Finally, Crossmodal Amplifier needs to be optimized, which includes the optimization of feature mapping and feature fusion parameters $\{w_{1},w_{2}\}$. To guide this process, as in Fig. 5, we define a total loss for Crossmodal Amplifier composed of two parts: the feature mapping loss $\mathcal{L}^{2D/3D}_{IBP}$, which aligns the mapped crossmodal features with the target encoder features, and the output consistency loss $\mathcal{L}^{2D/3D}_{output}$, which is used to adjust feature fusion parameters. The total loss function is
@@
The overall loss $\mathcal{L}_{TRD}$ is expressed as
$\mathcal{L}_{TRD}=(\mathcal{L}^{2D}_{D}+\mathcal{L}^{2D}_{CF}+\mathcal{L}^{2D}_{CA})+(\mathcal{L}^{3D}_{D}+\mathcal{L}^{3D}_{CF}+\mathcal{L}^{3D}_{CA})$ (9)
Since each part has independent optimizers and inputs, the hyper-parameters are set to 1.0.
Figure 6: Inference process of our proposed method. The feature evolution visualization highlights that anomalies in each modality are effectively captured. In resulting anomaly maps, darker gray regions indicate higher anomaly scores.
TABLE I: Anomaly Detection and Localization Results (%) on MVTec 3D-AD with the Best in Bold and the Second-best Underlined
Metrics Method Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
I-AUC BTF [7] 91.8 74.8 96.7 88.3 93.2 58.2 89.6 91.2 92.1 88.6 86.5
M3DM [8] 99.4 90.9 97.2 97.6 96.0 94.2 97.3 89.9 97.2 85.0 94.5
Shape-guided [32] 98.6 89.4 98.3 99.1 97.6 85.7 99.0 96.5 96.0 86.9 94.7
AST [5] 98.3 87.3 97.6 97.1 93.2 88.5 97.4 98.1 100 79.7 93.7
MMRD [9] 99.9 94.3 96.4 94.3 99.2 91.2 94.9 90.1 99.4 90.1 95.0
CFM [6] 98.8 87.5 98.4 99.2 99.7 92.4 96.4 94.9 97.9 95.0 96.0
TRD (Ours) 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
P-AUC BTF [7] 99.6 99.2 99.7 99.4 98.1 97.4 99.6 99.8 99.4 99.5 99.2
M3DM [8] 99.5 99.3 99.7 98.5 98.5 98.4 99.6 99.4 99.7 99.6 99.2
AST [5] - - - - - - - - - - 97.6
MMRD [9] - - - - - - - - - - 99.2
CFM [6] 99.7 99.2 99.9 97.2 98.7 99.3 99.8 99.9 99.8 99.8 99.3
TRD (Ours) 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
PRO BTF [7] 97.6 96.9 97.9 97.3 93.3 88.8 97.5 98.1 95.0 97.1 95.9
M3DM [8] 97.0 97.1 97.9 95.0 94.1 93.2 97.7 97.1 97.1 97.5 96.4
Shape-guided [32] 98.1 97.3 98.2 97.1 96.2 97.8 98.1 98.3 97.4 97.5 97.6
AST [5] 97.0 94.7 98.1 93.9 91.3 90.6 97.9 98.2 88.9 94.0 94.4
MMRD [9] 98.6 99.0 99.1 95.1 99.0 90.1 99.0 99.0 98.7 98.2 97.6
CFM [6] 98.0 96.6 98.2 94.7 95.9 96.7 98.2 98.3 97.6 98.2 97.2
TRD (Ours) 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
IV-D Inference
IV-D1 Anomaly Scoring
@@
IV-D2 Feature Evolution Analysis
To better understand the behavior of our TRD during inference, Fig. 6 visualizes the feature evolution process within TRD. Intuitively, Crossmodal Filter effectively supplements information from the other modality while suppressing anomaly signals that may interfere with anomaly-free feature reconstruction. In contrast, Crossmodal Amplifier allows anomaly information from one modality to be reflected in the other branch. Through the joint modulation of both components, each branch becomes aware of all modality-specific anomalies, resulting in more accurate and robust anomaly detection.
V Experiments
V-A Experimental Settings
Dataset
@@ V-A Experimental Settings
We conduct experiments on datasets MVTec 3D-AD [4] and Eyecandies [30]. MVTec 3D-AD is a 3D industrial anomaly detection dataset featuring 10 categories of industrial objects. The training set consists of 2,656 normal samples, each including an RGB image and corresponding 3D scan information, while the validation set contains 294 normal samples. The test set comprises 1,197 samples, including 948 anomalous samples and 249 normal samples, with pixel-level annotations provided for evaluation. For our experiments, we use the processed depth maps derived from the 3D scans as the 3D modality. Eyecandies is a synthetic dataset of 10 candy categories in an industrial conveyor scenario, providing RGB images, depth maps, and surface normal maps. It contains 10k training samples, 1k validation samples, and 4k test samples. In our experiments, we select normal maps as the 3D modality and chose images under consistent lighting conditions as the 2D modality.
Evaluation Metrics
@@
The evaluation metrics for anomaly detection include image-level area under the receiver operating characteristic curve (I-AUC) and average precision (I-AP). For anomaly localization, pixel-level AUC (P-AUC) and AP (P-AP) are used. We also employ the per-region overlap (PRO) metric calculated according to the official code [4] for anomaly localization, which treats anomaly regions of different size equally.
Implementation Details
@@
All RGB images, depth maps and normal maps are scaled to $\times 256$ 256. The training is performed on a single Nvidia GTX 3090 GPU, and a separate detection model is trained for each category. The network backbone is WideResNet-50 [38]. The batch size is 16, and the training lasts for 200 epochs. Each branch is optimized using a separate Adam optimizer with a learning rate of 0.005. In inference, the anomaly maps of both branches are smoothed using a Gaussian filter with $\sigma=4$.
TABLE II: Anomaly Detection and Localization Results (%) on Eyecandies with the Best in Bold and the Second-best Underlined
Metrics Methods
Candy
Cane
Chocolate
Cookie
Chocolate
Praline
Confetto
Gummy
Bear
Hazelnut
Truffle
Licorice
Sandwich
Lollipop
Marsh-
mallow
Peppermint
Candy
Average
I-AUC M3DM [8] 62.4 95.8 95.8 100 88.6 75.8 94.9 83.6 100 100 89.7
AST [5] 57.4 74.7 74.7 88.9 59.6 61.7 81.6 84.1 98.7 98.7 78.0
MMRD [9] 85.4 100 94.6 99.8 90.8 74.7 96.6 98.4 100 100 94.0
CFM [6] 68.0 93.1 95.2 88.0 86.5 78.2 91.7 84.0 99.8 96.2 88.1
TRD (Ours) 86.7 100 98.1 98.1 87.2 72.8 97.1 93.0 100 99.5 93.3
P-AUC M3DM [8] 97.4 98.7 96.2 99.8 96.6 94.1 97.3 98.4 99.6 98.5 97.7
AST [5] 76.3 96.0 91.1 96.9 78.8 83.7 91.8 92.4 98.3 96.8 90.2
MMRD [9] - - - - - - - - - - 98.3
CFM [6] 98.3 98.2 96.4 98.9 94.9 94.6 96.9 98.0 99.5 98.7 97.4
TRD (Ours) 99.5 99.0 99.0 99.5 95.7 94.6 99.6 99.0 99.7 99.5 98.5
PRO M3DM [8] 90.6 92.3 80.3 98.3 85.5 68.8 88.0 90.6 96.6 95.5 88.2
AST [5] 51.4 83.5 71.4 90.5 58.7 59.0 73.6 76.9 91.8 87.8 74.4
MMRD [9] 97.5 97.0 94.2 98.5 91.7 68.0 97.0 94.1 99.0 99.2 93.6
CFM [6] 94.2 90.2 83.1 88.0/96.5 87.5 76.2 79.1 91.3 93.9 94.9 88.7
TRD (Ours) 98.0 94.1 96.8 98.1 93.6 77.3 98.1 94.4 98.9 98.1 94.7
V-B Main Results
V-B1 Results on MVTec-3D AD
Table I shows the quantitative image-level anomaly detection results and pixel-level anomaly localization results on the MVTec 3D-AD dataset. We compare our method TRD with several state-of-the-art 3D anomaly detection methods.
@@ V-B1 Results on MVTec-3D AD
TRD consistently outperforms all other methods in both I-AUC and PRO metrics, achieving an average I-AUC of 96.2% and an average PRO of 97.7%. Notably, TRD demonstrates superior performance in detecting and localizing anomalies across a wide range of categories, including some challenging categories such as Cookie and Peach. A subset of qualitative results on MVTec 3D-AD is provided in Fig. 1, showing that our proposed TRD accurately localizes the anomaly regions.
Figure 7: Performance comparison of feature learning-based methods via radar chart.
TABLE III: Comparison of Feature Learning-based Methods on MVTec 3D-AD and Eyecandies datasets under the Same Hardware Setup (* indicates that MMRD has no reproducible code)
Method MVTec 3D-AD (%) Eyecandies (%)
Avg.
Frame
Rate (fps)
@@
TRD (Ours) 96.2 99.4 97.7 93.3 98.5 94.7 96.6 21.7
V-B2 Results on Eyecandies
We evaluate our method on the Eyecandies dataset, as shown in Table II. Our method TRD significantly outperforms all non-KD-based multimodal anomaly detection methods. While our TRD slightly lags behind MMRD [9] which uses both depth maps and surface normal maps as 3D modalities in anomaly detection metrics I-AUC, it consistently achieves superior performance 94.7% in anomaly localization metrics PRO.
V-B3 Comparison with Feature Learning-based Methods
We further compare TRD with feature learning-based methods including AST[5], MMRD[9], and CFM [6] on MVTec 3D-AD and Eyecandies. As in Fig. 7 and Table 7, TRD achieves the best average performance 96.6% and the highest frame rate 21.7 fps. Despite the multi-branch design, TRD remains lightweight due to its CNN-based architecture, outperforming other feature learning-based methods in efficiency.
TABLE IV: Anomaly Detection and Localization Results (%) of Different Modality Branches on MVTec 3D-AD
@@ V-B3 Comparison with Feature Learning-based Methods
2D+3D+CT (Ours) 96.2 99.4 97.7 98.8 45.5 87.5
Figure 8: Visualization of anomaly maps of different modality branches on MVTec 3D-AD.
Figure 9: Visualization of anomaly maps of different modality branches on Eyecandies.
V-B4 Singlemodal vs. Multimodal
We investigate the effectiveness of combining different modalities for AD through both quantitative and qualitative comparisons. We first compare singlemodal baselines (2D and 3D), their enhanced versions with Crossmodal Tuners (2D+CT and 3D+CT), and the full multimodal settings (2D+3D and 2D+3D+CT) in Table IV. Results on the MVTec 3D-AD dataset show that combining both modalities significantly improves performance. The I-AUC metric of 93.9% for using multiple modalities is higher than 86.8% for 2D only and 87.1% for 3D only, highlighting the complementary benefits of modality fusion.
To provide further insight, we visualize the anomaly maps of each branch under multimodal setup on MVTec 3D-AD and Eyecandies, as in Fig. 8 and Fig. 9. Some anomalies are hard to perceive in one modality but clearly visible in the other. Relying on a single modality may lead to missed detections, while our TRD designed for multimodal AD helps mitigate this issue through complementary information across modalities.
V-C Ablation Studies
TABLE V: Ablation Study Results (%) of Crossmodal Tuners on MVTec 3D-AD
@@ V-C Ablation Studies
Figure 10: Qualitative ablation results of Crossmodal Tuners on MVTec 3D-AD.
V-C1 Study on Crossmodal Tuners
We perform an ablation study to evaluate the impact of Crossmodal Tuners on anomaly detection performance, as shown in Table V. The results reveal that both CF and CA contribute to improved performance, with the combination of both achieving the best results. Specifically, adding CF to the baseline improves the I-AUC from 93.9% to 94.2%, and incorporating CA also boosts the performance, achieving an I-AUC of 95.6%. The highest performance is achieved when both CF and CA are used together, resulting in an I-AUC of 96.2%.
We also provide visualizations of anomaly maps before and after incorporating Crossmodal Tuners, as in Fig. 10. Without Crossmodal Tuners, anomalies that are less prominent in a specific modality goes undetected in the corresponding branch. However, after introducing Crossmodal Tuners, this issue is significantly mitigated, resulting in a noticeable reduction in the anomaly miss rate across all modality branches.
TABLE VI: Ablation Study Results (%) of Different Downsampling Sizes in Bottlneck Projection on MVTec 3D-AD
@@ V-C1 Study on Crossmodal Tuners
$\times 16$ 16 93.6 99.3 97.4 97.9 45.2 86.7
$\times 8$ 8 94.2 99.3 97.5 98.2 45.8 87.0
$\times 4$ 4 93.8 99.3 97.4 98.0 45.3 86.8
V-C2 Study on Bottlneck Projection in Crossmodal Filter
We investigate the effect of different downsampling sizes on the performance of the Crossmodal Filter, as shown in Table VI. When Bottleneck Projection reduces the feature size to $\times 8$ 8, the model achieves the best performance across multiple metrics. It is observed that excessively large downsampling size (e.g., $\times 4$ 4) leads to information loss, while small downsampling size (e.g., $\times 16$ 16) makes it difficult to filter out abnormal interference effectively. Therefore, we conclude that selecting an appropriate downsampling size is crucial for maintaining a balance between information preservation and interference removal.
TABLE VII: Ablation Study Results (%) of Different Channel Expansion Sizes in Inverted Bottlneck Projection on MVTec 3D-AD